OpenStudy (anonymous):

Why does the sum of the series (k!)^2/(2k)! from [1,infinity] converge?

OpenStudy (anonymous):

\[ \frac {(k!)^2} { (2 k)!}= \frac {(k!) } { (k+1)(k+2)(k+3)\cdots(k+k)}=\\ \frac 1{k+1}\frac 2{k+2}\frac 3{k+3}\cdots \frac k{k+k}\le\\ \frac 1 2\frac 1 2\frac 1 2\cdots \frac 1 2=\frac 1 {2^k} \] It converges by the comparison test.